OpenStudy (anonymous):

How do you find the sin(theta/2) given that cos(theta)=3/5, 0

OpenStudy (mathmale):

There's a so-called "half angle formula" for the sine; have you any familiarity with that?

OpenStudy (anonymous):

I have the formula

OpenStudy (anonymous):

Do i plug the cosine into the formula?

OpenStudy (mathmale):

Please see http://www.sosmath.com/trig/douangl/douangl.html From this reference I've gotten: \[\cos ^{2}\frac{ a }{ 2 }=\frac{ 1 }{ 2 }(1+\cos a).\]

OpenStudy (anonymous):

okay thanks

OpenStudy (mathmale):

Hold...still working on this.

OpenStudy (mathmale):

I should have given you the formula for the square of the SINE of (a/2). Sorry. But you can find that for yourself in the reference i've given you. Now to apply it:

OpenStudy (mathmale):

\[\sin ^{2}\frac{ \theta }{ 2 }=\frac{ 1 }{ 2 }(1-\cos \theta).\] You've been told that cos theta=3/5, and that theta is in Quadrant 1.

OpenStudy (mathmale):

Do the following: (1) substitute 3/5 for cos theta in the formula I've just given you. (2)Find the square root of both sides of this formula. this will give you an expression for sin (theta/2). Choose the positive root, not the negative one.

OpenStudy (mathmale):

Think: Why choose the positive root instead of the negative one?

OpenStudy (anonymous):

Because the first quadrant is positive?

OpenStudy (mathmale):

Yes, because the sine of theta/2 is the sine of a first-quadrant theta. All set?

OpenStudy (anonymous):

Yes. Thank you

OpenStudy (mathmale):

You're welcome. Hope to work with you again! Best wishes. :)